OpenStudy (anonymous):

How many moles of oxygen will occupy a volume of 2.5 liters at 1.2atm and 25 degrees C?

OpenStudy (anonymous):

PV=nRt [r=0.0821] derive a formula.. pv/RT=nRt/Rt n= pv/Rt p=1.2 atm v= 2.5 L t= 25 deg. C convert to Kelvin +273 = 298 K R= 0.0821 <constant n=(1.2)(2.5) / (0.0821)(298) n= 0.1226....... mol n= 0.12 mol [ accd. to significant figures] Does this help? ^v^

OpenStudy (anonymous):

thank you!!!!!!